OpenStudy (anonymous):

perpendicular to y=-2x +6 and containing the point (-2,4)

OpenStudy (anonymous):

Do you know the slope of your original line?

OpenStudy (anonymous):

No it part of my homework write equations for the following lines it says

OpenStudy (anonymous):

I understand. But the answer is right there. I was just wondering if you knew how to identify it.

OpenStudy (anonymous):

What how do you know I don't understand he told me to use this formula y=MX+b but I don't know how to set it up

OpenStudy (anonymous):

OK. Yes. That's the slope-intercept form of a line. y is your y coordinate m is your slope x is your x coordinate b is the y intercept (where the line crosses the y axis) So, now, looking at your original equation y=-2x +6 can you identify the slope?

OpenStudy (anonymous):

That's the problem I don't know I need an example then it will help

OpenStudy (shamil98):

y = mx + b m = slope b= y intercept y = -2x + 6 slope or m = -2/1 y intercept or b= 6

OpenStudy (shamil98):

Perpendicular lines have inverse slopes.

OpenStudy (shamil98):

The inverse of -2 /1 is 1/2

OpenStudy (anonymous):

cush needs to understand slopes before understanding inverse slopes. Giving away the answers doesn't help.

OpenStudy (shamil98):

Now to find the equation all you do is plug in those coordinates to find b y = mx + b 4 = -2(1/2) + b 4 = -1 + b 5 = b y = 1/2 x + 5

OpenStudy (shamil98):

cush said he needs an example.. that is an example..

OpenStudy (shamil98):

I've explained it.

OpenStudy (anonymous):

No, you gave him the answer. *I* explained it.

OpenStudy (anonymous):

Could you draw it because I don't get where you got the 1/2 and where the 4 or 5 came from

OpenStudy (shamil98):

(-2,4) (x,y) y = m(x) + b the slope of the original equation is -2/1 Perpendicular lines have inverse slopes ( the opposite ) , 1/2 in this case.. we have the slope and the (x,y) coordinates of the line but not the y-intercept at which you graph from. so we plug the slope and (x,y) coordinates into y =mx + b and solve for b. 4 = 1/2 (-2) + b 4 = -2/2 + b 4 = -1 + b add 1 to both sides 5 = b

OpenStudy (shamil98):

then input the values to get your equation.. y = 1/2 x + 5

OpenStudy (anonymous):

Omg I think I get it thank you can you tell me how to do the same thing with the same numbers and everything but with parrell lines

OpenStudy (shamil98):

Parallel lines have the same slope

OpenStudy (shamil98):

Could you give me your equation and coordinates?

OpenStudy (anonymous):

It has the same everything

OpenStudy (shamil98):

y=-2x +6 and containing the point (-2,4) parallel lines have the same slope so all you have to do is solve for b again :P

OpenStudy (shamil98):

I'll set the equation up for you 4 = -2(-2) + b

OpenStudy (anonymous):

Y=2x+4 and containing point 4,0 perpendicular I got y=-1/4+4 check if did it right

OpenStudy (shamil98):

4 = 4 + b b = 0 so that equation's y intercept is 0 y = 2x

OpenStudy (shamil98):

y= 2x + 4 y = 1/2 x + 4 is correct